When the air resistance is ignored, the horizontal range R of a projectile is given by R(θ) = v2 0 g sin 2θ where v0 is the constant initial velocity, g is the acceleration due to gravity, and θ is the angle of elevation or departure. Find the maximum range of projectile.


1
Expert's answer
2022-02-08T12:01:26-0500

Solution:

"R=\dfrac{v_{0}^{2} \sin 2 \theta}{g}"

This equation shows that for a given projectile velocity "\mathrm{v}_{0}, \mathrm{R}"  is maximum when "\sin 2 \theta"  is maximum, i.e. when "\theta_{0}=45^{\circ}" . The maximum horizontal range is, therefore

"{R}_{\mathrm{m}}=\dfrac{{v}_{0}{ }^{2}}{{~g}}"